OpenStudy (anonymous):

Let f(x)= (x^2 + 3x – 4) and g(x) = (x + 4).

OpenStudy (binary3i):

f=(x^2 +3x -4) =x^2 +4x -x -4 f =(x+4)(x-1) f.g= (x+4)(x-1)(x+4)=(x+4)^2(x-1) f/g=(x+4)(x-1)/(x+4)=(x-1)

OpenStudy (anonymous):

Thank you :)

OpenStudy (binary3i):

you'r welcome